# Asynchronous Federated Stochastic Optimization with Exact Averaging for Heterogeneous Local Objectives

### v1

@@ Abstract
Federated learning (FL) was recently proposed to securely train models with data held over multiple locations (“clients”) under the coordination of a central server. Two major challenges hindering the performance of FL algorithms are long training times caused by straggling clients and a decrease in training accuracy induced by non-iid local distributions (“client drift”). In this work we propose and analyze AREA, a new stochastic (sub)gradient algorithm that is robust to client drift and utilizes asynchronous communication to speed up convergence in the presence of stragglers. Moreover, AREA is, to the best of our knowledge, the first method that is both guaranteed to converge under arbitrarily long delays, and converges to an error neighborhood whose size depends only on the variance of the stochastic (sub)gradients used and thus is independent of both the heterogeneity between the local datasets and the length of client delays, without the use of delay-adaptive stepsizes. Our numerical results confirm our theoretical analysis and suggest that AREA outperforms state-of-the-art methods when local data are highly non-iid.
Keywords: Federated learning, asynchronous communication, non-iid data
@@ 1 Introduction
where the $i^{th}$ data-owning source (“client” $i\in\{1,...,n\}$) has private access to the function $f_{i}:\mathbb{R}^{p}\rightarrow\mathbb{R}$ representing the loss over its local data, $x\in\mathbb{R}^{p}$ is the vector of model parameters, and $\mathcal{X}\subseteq\mathbb{R}^{p}$ is a constraint set.
Throughout this work we will refer to $f(\cdot)$ and the functions $f_{i}(\cdot)$ for $i\in[n]$ as the global and local loss functions, respectively. A typical FL setup can be seen in Fig. 1; each client $i\in[n]$ optimizes the local loss $f_{i}(\cdot)$, obtains a model estimate $x_{i}\in\mathbb{R}^{p}$, and shares it with the server. The server waits until all clients have returned their estimates, calculates their average $\bar{x}=\frac{1}{n}\sum_{i=1}^{n}x_{i}\in\mathbb{R}^{p}$, and broadcasts it back to all clients. The clients receive $\bar{x}$ and use it as the starting point for the next optimization round.
Figure 1: Illustration of a federated learning (FL) architecture with heterogeneous devices on non-iid data.
Despite the ability to train a global model while maintaining the privacy of local data, FL faces two important challenges that hinder its performance: ($i$) the phenomenon of “client drift” [18] and ($ii$) the presence of slow clients (“stragglers”). In particular, when the local datasets are non-iid, clients may converge closer to the optima of their local losses and away from the optimum of the global loss, degrading the quality of the aggregated model (client drift). On the other hand, stragglers prolong and destabilize training while faster nodes remain idle. Asynchronous communication is an attractive solution to this issue; however, asynchronous algorithms are difficult to analyze, and the assumption of bounded delays may need to be adopted, which is impractical in real settings. This work aims to address all of these challenges, on which we elaborate further in the next two paragraphs.
In order to alleviate communication load and speed up convergence, the seminal FL algorithm FedAvg [28], also known as local stochastic gradient descent (SGD), and its various extensions [e.g., 18, 45] rely on the performance of multiple consecutive SGD steps at each client between communications with the server. However, multiple serial SGD steps may induce client drift under non-iid data. Moreover, the error caused by client drift is proportional to the heterogeneity between the local datasets [23, 45]. In the asynchronous setting, existing results [e.g., 51, 19, 43] suggest that asynchrony and longer client delays contribute to client drift even when a single SGD step is performed between updates, making it impossible to achieve exact convergence even when the true gradients are available (i.e., in the absence of stochastic gradient noise).
In addition to client drift, we also consider in this work the challenge of stragglers in FL, and the lack of theoretical guarantees for asynchronous methods without impractical bounds on the client delays. Although the poor scalability of synchronous methods and the presence of straggling nodes have led to a considerable interest in FL methods utilizing asynchronous updates [49], the theoretical study of asynchronous algorithms is particularly challenging, as evident from the various frameworks proposed for this purpose over the years [e.g., 26, 14, 13]. Existing analyses typically rely on deterministic bounds on client delays, a setting known as “partially asynchronous” [4]. Partially asynchronous algorithms are not guaranteed to converge under arbitrarily long delays, while algorithms providing such guarantees are known as “totally asynchronous”. A framework for analyzing totally asynchronous iterations with applications in FL was recently proposed in [13]; however, convergence rate guarantees are limited to the case where delays grow with a slower rate than time itself.
Contributions. In this paper we propose and analyze AsynchRonous Exact Averaging (AREA), a new stochastic (sub)gradient method that addresses the challenges of straggling clients and client drift in FL. Specifically, AREA utilizes asynchronous communication to avoid overly long training times caused by stragglers, and employs correction terms stored in memory to counteract the deviation from the optimum of the global loss induced by non-iid data. In addition, our theoretical results do not require the impractical assumption of bounded client delays; namely, AREA is a totally asynchronous method. A significant advantage of AREA over existing works, is that it provably converges to an error neighborhood that is independent of the length of client delays, without the use of delay-adaptive stepsizes; namely a constant stepsize can be utilized throughout the optimization process, which can furthermore be tuned without any prior knowledge of the system delays. To the best of our knowledge, AREA is the first totally asynchronous method that is also robust to client drift and enjoys such theoretical guarantees.
Our contributions can be summarized as follows:
@@
  • We propose AREA, a new stochastic (sub)gradient method for FL with non-iid data, which utilizes asynchronous updates to address the issues of slow and unstable convergence caused by straggling clients.
  • We provide rigorous theoretical results on the convergence of AREA in both convex and strongly convex settings. Specifically, for strongly convex and smooth functions we show that AREA with constant and delay-independent stepsize converges with linear rate to a neighborhood of the optimal solution of Problem (1), whose size depends only on the variance of the stochastic gradients and is thus independent of both the client delays and the heterogeneity between the local datasets. For constrained problems and convex and nonsmooth functions, we recover the convergence rate of the centralized stochastic projected subgradient method [7].
@@
  • Our analysis does not require deterministic bounds on the client delays, making AREA a totally asynchronous algorithm [4]. Compared with the limited subset of works in the totally asynchronous setting [15, 13], we are able to establish error bounds that are independent of client delays, without the use of delay-adaptive stepsizes.
@@
  • For our theoretical analysis we prove an extended version (Lemma 20 in Appendix A) of the well-known result of Robbins–Siegmund [see 36, Lemma 11, Ch. 2], which is often used in the literature to derive the almost sure convergence of random sequences. This extended result can be used to construct suitable Lyapunov functions tracking the progress of asynchronous iterations and could be of independent interest in the probabilistic study of algorithms.
@@ 2 Related Work
Non-iid data/heterogeneous local objectives. Several methods have been proposed to tackle client drift in synchronous FL [22, 34, 18, 45]; however, none eliminate it completely [31]. The synchronous methods FedLin [31] and Scaffnew [30] employ gradient correction terms to steer convergence towards the correct optimum. Moreover, FedLin allows for heterogeneous amounts of local computation to mitigate the effect of straggling nodes but requires the two-way communication of both parameter and gradient vectors. Moreover, the aggregation of the latter is executed synchronously. Hence, all aforementioned methods may suffer from slow convergence in the presence of stragglers and may not scale favorably as the number of clients increases. In this work, we adopt asynchronous communication to enhance scalability and speed up convergence. To the best of our knowledge, only a small subset of asynchronous federated methods is able to mitigate client drift [3, 15, 47, 50]; however the error bounds provided in all of these works grow with data heterogeneity, even in the case where a single SGD step is executed locally. Moreover, with the exception of [15], all aforementioned works are limited to the partially asynchronous setting. The algorithm MIFA [15], the closest work to ours, can tolerate growing client delays over time, but the authors do not establish the independence of MIFA’s convergence neighborhood from the average and maximum delays. In contrast, we are able to theoretically demonstrate this independence for AREA.
Asynchronous federated learning. Existing approaches for addressing staleness and client disparities in FL include (i) semi-synchronous strategies such as adjusting the workload at each client [8, 24] or forcing clients to communicate inexact updates upon server request [52, 21] and (ii) fully asynchronous techniques such as employing custom aggregation policies [33, 14, 48, 46] and delay adaptive stepsizes [38, 9]. However, all of these methods are sensitive to client drift, and all are partially asynchronous, with the exception of the works employing delay adaptive stepsizes [38, 9] and those relying on strong assumptions such as client responsiveness upon server request [52, 21]. Conversely, AREA is totally asynchronous, robust to client drift and utilizes delay-independent stepsizes.
Asynchronous SGD. Asynchronous SGD and its variants are the cornerstones of large-scale parallel and distributed optimization [44, 10, 37, 35, 42]. The convergence properties of parallel and distributed SGD, mini-batching, and local SGD with asynchronous updates have been studied extensively in the literature [e.g., 1, 40, 41, 12, 53, 25, 2, 39, 19]. However, existing analyses typically enforce deterministic bounds on the maximum or average client delays, which may also be required to be known a priori to tune the stepsize. The authors of [13] employ their framework to study the asynchronous version of FedAvg under total asynchrony, however their analysis is limited to the iid case and requires a delay-adaptive stepsize rule to guarantee convergence. Conversely, we are able to handle total asynchrony with constant stepsize that is delay-independent. Finally, unlike AREA, asynchronous SGD and its variants are known to be susceptible to client drift.
@@ 3 AREA: An asynchronous method for federated stochastic optimization
In this section we propose a new stochastic (sub)gradient algorithm, AREA, to solve Problem (1) in a decentralized manner while addressing the slow and unstable convergence caused by straggling clients. We consider the case where client $i\in[n]$ can obtain only a stochastic approximation of the true (sub)gradient of $f_{i}(\cdot)$, which satisfies the following standard assumption.
@@
Let $d_{i}(x)$ be the (sub)gradient of $f_{i}(\cdot)$ at $x\in\mathbb{R}^{p}$, and let $g_{i}(x)$ be a stochastic approximation of $d_{i}(x)$. Then for all $i\in[n]$ and some $\sigma>0$, the following hold: $\mathbb{E}[g_{i}(x)-d_{i}(x)|x]=0$ and $\mathbb{E}[\|g_{i}(x)-d_{i}(x)\|^{2}|x]\leq\sigma^{2}$, where the expectation is taken over all possible realizations of $g_{i}(x)$.
We note that stronger results can be achieved with the true (sub)gradient, which is discussed in Section 4.2.
@@
Initialization: initialize model $x_{s}\in\mathcal{X}$, aggregator $u_{s}=0$, count $k=0$ and stepsize $\alpha_{0}>0$; broadcast $x_{s}$ and $\alpha_{0}$ to all clients $i\in[n]$
@@
         update $k\leftarrow k+1$ and construct $\alpha_{k}$
         send $x_{s}$ and $\alpha_{k}$ to client $i$
     end if
     if the aggregation criterion satisfied then
         update $x_{s}\leftarrow x_{s}+u_{s}$, $u_{s}\leftarrow 0$ and $k\leftarrow k+1$
     end if
end while
@@
The optimization process of AREA at the server is summarized in Algorithm 1. The server maintains the global variable $x_{s}\in\mathcal{X}$, an aggregator variable $u_{s}\in\mathbb{R}^{p}$ tracking the average variables across all clients, and a counter $k\in\mathbb{N}_{0}$ of the number of client updates the server has received. During the initialization phase, the server sets the aggregator $u_{s}$ and the counter $k$ equal to zero and initializes $x_{s}\in\mathcal{X}$ and a stepsize sequence $\{\alpha_{k}\}$. The server then broadcasts $x_{s}$ and $\alpha_{0}$ to all clients. Upon receipt of an update $m_{i}$ from client $i$, the server increases the counter $k$ by $1$ and sends $\alpha_{k}$ to client $i$ with the current version of $x_{s}$. Moreover, the server adds the quantity $m_{i}/n$ to the aggregator $u_{s}$. Next, the server checks whether a certain aggregation criterion is satisfied; if it is, the server updates $x_{s}$ by adding to its current value the contents of $u_{s}$, resets the value of $u_{s}$ to zero, and increases the counter $k$ by $1$. These steps are repeated until convergence is reached and the server broadcasts a stopping signal to all clients.
@@
The aggregation criterion can be selected on a case-by-case basis to meet application demands; possible options include updating the global variable every $\Delta\in\mathbb{N}$ client communications (e.g., [33]), after fixed time intervals (e.g., [14]), or even randomly. We note that AREA is a secure aggregation (SecAgg) friendly method and can provide protection against honest-but-curious server attacks [6, 33].
@@
Initialization: receive global model $x_{s}$ and set $y_{i}=x_{s}$
@@
     receive $x_{s}$ and $\alpha_{k}$ from server
     calculate stochastic (sub)gradient $g_{i}(x_{s})$ of $f_{i}(\cdot)$ at $x_{s}$
     update $x_{i}\leftarrow P_{\mathcal{X}}[x_{s}-\alpha_{k}g_{i}(x_{s})]$
@@
end while
The process of AREA at client $i$ for $i\in[n]$ is shown in Algorithm 2. We note that Algorithms 1 and 2 run in parallel and transfer information to each other over some communication protocol. Each client $i\in[n]$ maintains two local variables $x_{i},y_{i}\in\mathcal{X}$, storing the most recent and the immediately previous local estimate of the global variable, respectively. We note that $y_{i}$ requires storage between communications with the server but not $x_{i}$. During the initialization phase, the clients receive $\alpha_{0}$ and $x_{s}$ and set $y_{i}=x_{s}$ for all $i\in[n]$. The following steps are then repeated by all clients until a stopping signal from the server is received: upon receipt of a global model $x_{s}$ and a stepsize value $\alpha_{k}$ from the server, client $i\in[n]$ computes at its own pace a stochastic (sub)gradient approximation $g_{i}(x_{s})$ of $f_{i}(\cdot)$ at $x_{s}$ and takes a stochastic (sub)gradient step. The result of this operation is projected to $\mathcal{X}$ and stored in $x_{i}$ (we note that if $\mathcal{X}=\mathbb{R}^{p}$, the projection operator is the identity matrix). Client $i$ then constructs the message $m_{i}=x_{i}-y_{i}$ and sends it to the server along with a request for the most recent version of $x_{s}$. Once client $i$ has communicated $m_{i}$ to the server, it stores the current value of $x_{i}$ in $y_{i}$ and releases the memory for $x_{i}$.
@@
The communication round-based FedLaAvg (Algorithm 2 in [50]) with a single local SGD step at each client follows a similar procedure to AREA; however, FedLaAvg utilizes a different aggregation policy where only the updates from those clients that have not communicated with the server for the longest time are used for the global update; this may add bias in favor of some clients. Moreover, FedLaAvg provides guarantees only in the partially asynchronous setting.

### v2

# Asynchronous Federated Stochastic Optimization for Heterogeneous Objectives Under Arbitrary Delays

@@ Abstract
Federated learning (FL) was recently proposed to securely train models with data held over multiple locations (“clients”) under the coordination of a central server. Two major challenges hindering the performance of FL algorithms are long training times caused by straggling clients, and a decline in model accuracy under non-iid local data distributions (“client drift”). In this work, we propose and analyze AsynchRonous Exact Averaging (AREA), a new stochastic (sub)gradient algorithm that utilizes asynchronous communication to speed up convergence and enhance scalability, and employs client memory to correct the client drift caused by variations in client update frequencies. Moreover, AREA is, to the best of our knowledge, the first method that is guaranteed to converge under arbitrarily long delays, without the use of delay-adaptive stepsizes, and (i) for strongly convex, smooth functions, asymptotically converges to an error neighborhood whose size depends only on the variance of the stochastic gradients used with respect to the number of iterations, and (ii) for convex, non-smooth functions, matches the convergence rate of the centralized stochastic subgradient method up to a constant factor, which depends on the average of the individual client update frequencies instead of their minimum (or maximum). Our numerical results validate our theoretical analysis and indicate AREA outperforms state-of-the-art methods when local data are highly non-iid, especially as the number of clients grows.
@@ 1 Introduction
where the $i^{th}$ data-owning source (client $i\in\{1,...,n\}$) has private access to the function $f_{i}:\mathbb{R}^{p}\rightarrow\mathbb{R}$ representing the loss over its local data, $x\in\mathbb{R}^{p}$ is the vector of model parameters, and $\mathcal{X}\subseteq\mathbb{R}^{p}$ is a constraint set. Throughout this work, we will refer to $f(\cdot)$ and the functions $f_{i}(\cdot)$ for $i\in\{1,...,n\}$ as the global and local loss functions, respectively.
At each round of a typical FL algorithm, clients use their local data to independently optimize the model parameters and share their estimates with the server. The server then produces the global model via an aggregation process (e.g., weighted averaging). Although FL enables learning rich models from decentralized, private data that may otherwise not be available, it may suffer from long training times due to the presence of slow clients (“stragglers”). Asynchronous communication can speed up convergence in the presence of stragglers by preventing fast clients from remaining idle, in addition to removing the need for synchronization which may be infeasible in large networks.
Asynchronous iterations, however, are difficult to analyze theoretically, as evidenced from the various frameworks proposed for this purpose over the years [27, 16, 15, 14]. The assumption of bounded client delays is frequently adopted to guarantee convergence (e.g., [26, 3, 44, 20, 34]), a setting known in the literature as “partially asynchronous” [4]; however, this assumption is largely impractical in FL, where clients behave inconsistenly over time. A more realistic model of asynchronous communication over networks is “total asynchrony” [4], i.e. arbitrarily long delays between iterations. In the FL setting, totally asynchronous iterations are typically analyzed under delay-adaptive stepsizes [40, 20, 51, 30, 15, 51], which may require small learning rates to accommodate long delays and hence may result in slow convergence. Unlike existing approaches, we adopt both unbounded delays and delay-independent stepsizes, and recover explicit convergence rates with respect to the number of iterations for non-iid local data distributions and (i) strongly convex and smooth functions and (ii) convex and non-smooth functions. To the best of our knowledge, the only work providing convergence guarantees in the non-iid data setting, under strong convexity and smoothness and linearly growing, unbounded delays is [17]; however, the resulting rates grow with the maximum delay between iterations. Conversely, we adopt a fully probabilistic framework (i.e., we make assumptions only on the statistics of client delays) and recover rates depending only on the average client update frequency, rendering our method less sensitive to fast or slow clients. In the case of convex and non-smooth functions and non-iid local data, the algorithmic framework in [14] does not impose restrictions on the delays or the stepsize; however, the resulting rates depend on the maximum client sampling probability. In contrast to [14], we allow clients to update the server at their own pace, and recover rates depending only on the average probabilities.
Another challenge hampering the performance of FL, addressed in this work, is the phenomenon of “client drift” [19, 48]. To reduce communication overhead, the seminal FL algorithm FedAvg [29], also known as local stochastic gradient descent (SGD), executes multiple consecutive SGD steps at each client before aggregations at the server. Under non-iid local data, however, performing multiple SGD steps may cause clients to converge close to the minima of their local losses instead of the minimum of the global loss, degrading the quality of the aggregated model (client drift) [24]. Asynchronous communication may exacerbate this deviation, with frequently updating clients being over-represented and slow clients underrepresented in the global model. Existing results for asynchronous FL (e.g., [54, 20, 46]) indicate that the convergence error grows with the product of client delays and the heterogeneity between the local datasets, even when a single SGD step is executed locally. To eliminate the client drift caused by variations in client update frequencies, we propose an error compensation scheme that utilizes client memory for the storage of correction terms. The effectiveness of our approach is validated by our theoretical results, which are independent of the heterogeneity between the local datasets when one SGD step is executed locally. A number of existing methods employ similar mechanisms for the same purpose [3, 17, 50, 53], but unlike this work, they do not confirm the elimination of client drift theoretically.
Contributions. We propose and analyze AsynchRonous Exact Averaging (AREA), a new stochastic (sub)gradient method that addresses the challenges of stragglers and scalability in FL by adopting asynchronous communication. To correct the client drift caused by heterogeneous client update frequencies, AREA utilizes a memory-based error compensation scheme. Moreover, our theoretical results do not require the unrealistic assumption of bounded delays to guarantee convergence, making AREA a totally asynchronous method. Our contributions can be summarized as follows:
@@
  • We propose AREA, a new stochastic (sub)gradient method for FL with non-iid data, which utilizes asynchronous updates to address the issues of limited scalability and slow convergence caused by straggling clients.
  • AREA employs a memory-based compensation mechanism to correct the client drift caused by heterogeneous client update frequencies. Although other works utilize similar schemes [3, 17, 50, 53], we are the first to provide tight bounds which theoretically confirm the elimination of client drift for strongly convex and smooth functions. For convex and non-smooth functions and non-iid data, we are the first to report convergence rates matching the centralized projected stochastic subgradient method [7], up to a constant factor that depends only on the average client update frequency. Conversely, the only known rate in this setting grows with the maximum client update frequency [14], which may be significantly larger. Moreover, [14] adopts a biased sampling scheme for handling non-iid data, while we allow clients to update the server at their own pace. For a detailed comparison between our results and the state-of-the-art in the settings we study, see Tables 1 and 2 and Section 4.1.
@@
  • Our analysis does not require bounded client delays to guarantee convergence, making AREA a totally asynchronous algorithm [4], without the use of delay-adaptive stepsizes.
@@
  • For our theoretical analysis we prove an extended version (Lemma 20 in Appendix A.1) of the well-known result of Robbins–Siegmund (see [38], Lemma 11, Ch. 2), which is often used to derive the almost sure convergence of random sequences. This extended result can be used to construct suitable Lyapunov functions for tracking the progress of asynchronous iterations, and could be of independent interest in the probabilistic study of algorithms.
@@
Table 1: Summary of asynchronous FL methods analyzed under strong convexity and smoothness. We refer to the asynchronous version of SGD as AsyncSGD. The works [43, 30, 15] adopt delay-adaptive stepsizes, while [17] assumes unbounded delays restricted to linear growth. For all methods, we consider the case where a single SGD is perfomed locally at each client before communicating with the server. We note that, although [3, 17] may correct the client drift due to asynchrony in practice, this is not verified theoretically; see Section 4.1 for more details.
Reference Non-iid data Unbounded delays No client drift
AsyncSGD [30] N/A
AsyncSGD [15] N/A
AsyncSGD [43] N/A
AsyncSGD [8]
AsyncCommSGD [3]
MIFA [17]
AREA (this work)
Table 2: Summary of asynchronous FL methods analyzed under convex and non-smooth functions $f_{i}(\cdot)$ in Problem (1). We refer to the asynchronous version of SGD as AsyncSGD.
Reference Non-iid data No client sampling
AsyncSGD [30]
AsyncSGD [14]
AREA (this work)
@@ 2 Related Work
Client drift in FL. Several methods have been proposed for tackling the client drift due to multiple local SGD steps in synchronous FL [23, 36, 19, 48]; however, none eliminate it completely [32]. FedLin [32] and Scaffnew [31] use gradient correcting terms to steer convergence towards the correct minimum, but require synchronous communication which limits their practical applicability. A small subset of asynchronous FL methods employ memory to remedy the client drift due to heterogeneous client update frequencies [3, 17, 50, 53]; however, they all require bounded client delays, with the exception of [17] which adopts linearly growing unbounded delays.
Asynchronous FL. Existing approaches for addressing stale updates and heterogeneous client update frequencies in asynchronous FL include (i) semi-synchronous strategies, such as adjusting the workload at each client [8, 25], or forcing clients to communicate inexact updates upon server request [55, 22] and (ii) fully asynchronous techniques, such as custom aggregation policies [35, 16, 52, 49] and delay adaptive stepsizes [40, 10, 20, 51, 30]. However, all methods listed above suffer from client drift due to variations in client update frequencies under non-iid data. Moreover, with the exception of the works that employ delay-adaptive stepsizes, all aforementioned works require bounded client delays to guarantee convergence. As we previously noted, methods utilizing delay-adaptive stepsizes may adopt small learning rates to compensate for extended delays which may slow down convergence.
AsynchronousSGD. Asynchronous SGD and its variants are the cornerstones of large-scale parallel and distributed optimization [47, 11, 39, 37, 45]. The convergence properties of parallel and distributed SGD [13, 1, 20, 43, 56, 26, 41, 2], and local SGD [42, 8, 30] with asynchronous or delayed updates have been studied extensively in the literature. For a detailed comparison between our results and existing results on asynchronous SGD in the settings we study see Tables 1 and 2.
@@ 3 AREA: An Asynchronous Method for Federated Stochastic Optimization
We propose a new asynchronous stochastic (sub)gradient algorithm, AREA, to solve Problem (1) in a distributed manner under arbitrary delays. We assume that client $i\in[n]$ can obtain only a stochastic approximation of the true (sub)gradient of $f_{i}(\cdot)$ (e.g., stochatic (sub)gradient, mini-batch), satisfying the following standard assumption.
@@
Let $d_{i}(x)$ be the (sub)gradient of $f_{i}(\cdot)$ at $x\in\mathbb{R}^{p}$, and let $g_{i}(x)$ be a stochastic approximation of $d_{i}(x)$. Then for all $i\in[n]$ and some $\sigma>0$, we have: $\mathbb{E}[g_{i}(x)-d_{i}(x)|x]=0$ and $\mathbb{E}[\|g_{i}(x)-d_{i}(x)\|^{2}|x]\leq\sigma^{2}$, where the expectation is taken over all possible realizations of $g_{i}(x)$.
@@
Initialization: initialize model $x_{s}\in\mathcal{X}\subseteq\mathbb{R}^{p}$, aggregator $u_{s}=0\in\mathbb{R}^{p}$, stepsize $\alpha>0$ and stepsize update function $h:\mathbb{R}\rightarrow\mathbb{R}$; broadcast $x_{s}$ and $\alpha$ to all clients $i\in[n]$
@@
         update $\alpha\leftarrow h(\alpha)$ and send $x_{s}$ and $\alpha$ to client $i$
     if the aggregation criterion is satisfied then
         update $x_{s}\leftarrow x_{s}+u_{s}$, $u_{s}\leftarrow 0$
@@
The procedure of AREA at the server is summarized in Algorithm 1. The server maintains the global model $x_{s}\in\mathcal{X}$, an aggregator $u_{s}\in\mathbb{R}^{p}$, and a stepsize update function $h:\mathbb{R}\rightarrow\mathbb{R}$ (e.g., $h(x)=x$ for constant stepsize). At initialization, the server sets $u_{s}=0\in\mathbb{R}^{p}$, and initializes $x_{s}\in\mathcal{X}$ and the stepsize $\alpha$ and broadcasts them to all clients. Whenever an update $m_{i}\in\mathbb{R}^{p}$ arrives from some client $i\in[n]$, the server adds $m_{i}/n$ to $u_{s}$, constructs the new stepsize $h(\alpha)$ and sends it to client $i$ with $x_{s}$. In addition, the server is triggered whenever a pre-defined aggregation criterion is satisfied; the server then updates $x_{s}$ by adding to it the contents of $u_{s}$ and resets the value of $u_{s}$ to zero. These steps are repeated until convergence is reached and the server broadcasts a stopping signal to all clients.
@@
The aggregation criterion can be selected on a case-by-case basis to meet application demands. Possible options include aggregating every $\Delta\in\mathbb{N}$ client communications (e.g., [35]), after fixed time intervals (e.g., [16]), or even randomly. We note that it is possible to incorporate secure aggregation (SecAgg) to AREA for protection against honest-but-curious server attacks [6, 35].
@@
Initialization: receive global model $x_{s}$ and set $y_{i}=x_{i}=x_{s}$
@@
     receive $x_{s}$ and $\alpha$ from server and update $x_{i}\leftarrow P_{\mathcal{X}}[x_{s}-\alpha g_{i}(x_{s})]$
@@
The process of AREA at client $i\in[n]$ is shown in Algorithm 2. Client $i\in[n]$ maintains two variables $x_{i}\in\mathcal{X}$ and $y_{i}\in\mathcal{X}$, storing the two most recent local estimates of the global model. During the initialization phase, clients receive $\alpha$ and $x_{s}$ and set $y_{i}=x_{i}=x_{s}$. The following steps are then repeated by all clients until a stopping signal from the server is received: whenever a global model $x_{s}$ and a stepsize value $\alpha$ are received from the server, client $i\in[n]$ computes at its own pace a stochastic (sub)gradient approximation $g_{i}(x_{s})$ of $f_{i}(\cdot)$ and takes a stochastic (sub)gradient step of length $\alpha_{k}$. The result of this operation is projected to $\mathcal{X}$ and stored in $x_{i}$. Next, client $i$ constructs the message $m_{i}=x_{i}-y_{i}$ and sends it to the server along with a request for the most recent version of $x_{s}$. Once client $i$ has communicated $m_{i}$ to the server, it stores the current value of $x_{i}$ in $y_{i}$.
@@
The communication round-based FedLaAvg (Algorithm 2 in [53]) with a single local SGD step at each client follows a similar procedure to AREA; however, FedLaAvg utilizes a stricter aggregation policy, where only the updates from clients who have not recently communicated with the server are used in the aggregation and the rest are discarded.